How do u find the resultant vector's magnitude as well as anlge using 5.0 m/s, 45 deg + 2.0 m/s, 180 deg
the ansner is = 3.85m/s, 66.5 deg but i would like to know how to do it and the formula you have used

Answer by Alan3354(69443)

How do u find the resultant vector's magnitude as well as anlge using 5.0 m/s, 45 deg + 2.0 m/s, 180 deg
the ansner is = 3.85m/s, 66.5 deg but i would like to know how to do it and the formula you have used
------------------------------
Resolve the 2 vectors into x and y components: i = x dir, j = y dir
5.0 m/s, 45 deg --> 5*sin(45)i +5*cos(45)j
2.0 m/s, 180 deg --> -2i
Add the components
= 1.5355i + 3.5535j (the resultant vector)
Magnitude = sqrt(i^2 + j^2) = sqrt(2.3579 + 12.5) = 3.85 m/sec
direction = arctan(j/i) = arctan(2.3025)
direction = 66.5 degs
------------------------
There's a better method, if you know vectors.